Jack bought 3 slices of cheese pizza and 4 slices of mushroom pizza for a total cost of $12.50. Grace bought 3 slices of cheese pizza and 2 slices of mushroom pizza for a total cost of $8.50. What is the cost of one slice of mushroom pizza?

Found 2 solutions by FrankM, stanbon:
Answer by FrankM(1040)

3C + 4M = 1250
3C + 2M = 850
subtract,
2M = 400
M = 200
3C +2(200) = 850
3C + 400 = 850
3C = 450
C = 150
mushroom slices are $2.00, plain cheese $1.50

Answer by stanbon(75887)

Jack bought 3 slices of cheese pizza and 4 slices of mushroom pizza for a total cost of $12.50. Grace bought 3 slices of cheese pizza and 2 slices of mushroom pizza for a total cost of $8.50. What is the cost of one slice of mushroom pizza?
--------------
Equations:
3c + 4m = 12.50
3c + 2m = 8.50
-----------------------
Subtract and solve for "m"::
2m = $4
---
Ans: m = $2.00
---------------------
Cheers,
Stan H.
------------